OpenStudy (highschoolmom2010):

Use the given circumference to find the surface area of each spherical object. Round your answer to the nearest whole number. pincushion with C=8 cm

OpenStudy (ranga):

Circumference = 2(pi)r = 8 cm. r = 8 / (2pi) Surface area of sphere = 4(pi)r^2 = 4(pi) * 64 / (4pi^2) = 64 / pi = ? cm^2

OpenStudy (highschoolmom2010):

\[C=2\pi r\] \[8=2\pi r\] \[\frac{ 8 }{ 2\pi }=\frac{ 2\pi r }{2 \pi }\] \[1.27=r\] \[SA=4\pi r^2\] \[SA=4 \pi (1.27)^2\] \[SA=20.27\] \[SA \approx 20 cm ^3\]

OpenStudy (highschoolmom2010):

sorry i meant 20 cm ^3

OpenStudy (ranga):

Yes, area rounded to the nearest whole number is: 20 cm^2

OpenStudy (highschoolmom2010):

hooray :D

OpenStudy (highschoolmom2010):

ty @ranga

OpenStudy (ranga):

yw.